OpenStudy (anonymous):

fx=x sqrt of(6-x) looking for f'??

OpenStudy (yttrium):

Do you know product rule?

OpenStudy (anonymous):

yes, that is what I was thinking.

OpenStudy (yttrium):

Well you can use that.

OpenStudy (anonymous):

Would you convert the sqrt to 6-x ^1/2??

OpenStudy (anonymous):

And use the power rule?

OpenStudy (yttrium):

Well conversion and power rule will also do. Look at this: By converting the whole expression to a single power: \[f(x) = \sqrt{6x^2 - x^3} = (6x^2 - x^3)^{\frac{ 1 }{ 2 }}\]

OpenStudy (yttrium):

^now you can use power rule there

OpenStudy (anonymous):

Outstanding, let me work on this, Thank you!!

OpenStudy (yttrium):

Okay. Good luck. :)

OpenStudy (anonymous):

Help again, how did you get those values. Can you break it down for me.

OpenStudy (yttrium):

What values are you talking about?

OpenStudy (anonymous):

\[fx=x \sqrt{6-x}\] This is what I started with. Did you multiply the x into the radical. Don't you need to multiply radical w/ radical? not sure of this part.

OpenStudy (anonymous):

\[x ^{1}*6-x ^{1/2}?\]

OpenStudy (yttrium):

aah okay. let have one basic example \[\sqrt{x^2} = x \] so if you wanna convert x into radical form, you'll make it squared so in that case: \[f(x) = \sqrt {x^2 (6-x)}\] hence \[f(x) = \sqrt { 6x^2 - x^3}\]

OpenStudy (anonymous):

You are the bomb. Thanks. How does the medals thing work here?